OpenStudy (anonymous):

What is the vertex of a parabola whose equation is y = (x - 3)2? (-3, 0) (0, -3) (0, 3) (3, 0)

OpenStudy (marco26):

y-h=(x-k)^2 (h,k) is the vertex

OpenStudy (anonymous):

y = x^2 - 3 - 1 y = x^2- 3 y =x + 2 - 9 vertex: (0 , -3)

OpenStudy (anonymous):

Vertex of a quadratic function of the form y=a(x-h)^2 +k is at (h,k). This problem follows the pattern, so the vertex is also the x intercept at (3,0).